OpenStudy (anonymous):

Help[ pleaseeeeeeee Find the area of a rectangular basketball court if the length is \sqrt {8839} ft. and the width is \sqrt {2500} ft. Round your answer to the nearest square foot

OpenStudy (aaronq):

A=l*w

OpenStudy (anonymous):

8839*2500=22097500

OpenStudy (aaronq):

you didn't write the squared roots \[\sqrt{8839}*\sqrt{2500}\]

OpenStudy (anonymous):

I don't have it on my keyboard

OpenStudy (aaronq):

no worries, you still need to evaluate it correctly though

OpenStudy (anonymous):

how did I not do it correct

OpenStudy (aaronq):

you didn't take their square roots before multiplying them together

OpenStudy (anonymous):

94.01595609256974

OpenStudy (anonymous):

that's square for 8839

OpenStudy (anonymous):

and the other is 50

OpenStudy (aaronq):

squared root, you mean :P

OpenStudy (aaronq):

so just multiply them together now

OpenStudy (anonymous):

yea sorry

OpenStudy (aaronq):

lol no worries

OpenStudy (anonymous):

I got 4700.797805 ???

OpenStudy (aaronq):

yep, now, round your answer to the nearest square foot

OpenStudy (anonymous):

how do u do that

OpenStudy (anonymous):

so is it 5 ?

OpenStudy (aaronq):

4700.79, to the nearest integer, so 4700 b/c 0.79 is closer to 1 than to 0